# Retrieval augmented text-to-SQL generation for epidemiological question answering using electronic health records

### v1

@@ 1 Introduction
  • We show that retrieval augmented generation (RAG) significantly improve performance compared with static instruction prompting, as confirmed by extensive benchmarking with top-tier large language models (LLMs).
  • We share our dataset, code, and prompts111upon acceptance of the paper to foster reproducibility and catalyse a community-driven effort towards advancing this research area.
@@ 2 The Dataset
# of question/SQL pairs (all)
306
# of different tables used (all)
13
# of different columns used (all)
44
# logical conditions/query
6.4 (6.7)
# nesting levels/query
1.5 (1.1)
# tables/query
2.7 (0.9)
# columns/query
6.3 (4.7)
# medical entities/query
2.0 (4.1)
Question length [char]/query
91.7 (81.2)
SQL query length [char]/query
796.4 (448.5)
Table 1: Summary statistics of the dataset. For sample statistics, average and standard deviation (in bracket) are reported.
@@ 4.1 Experimental setup
Database and evaluation. The evaluation data reported are obtained using the DE-SynPUF dataset SynPUF (2010), which is a synthetic dataset including 6.8 million beneficiary records, 112 million claims records, and 111 million prescription drug events records Gonzales et al. (2023). The same analysis could be applied to any database conforming to the OMOP-CDM, thus potentially allowing access to 2.1 billion patient records. For evaluation, we manually compare the data retrieved via the reference and generated query. A generated query is marked as correct if it retrieves data enabling an answer that aligns with the reference query’s answer (within a 10% tolerance), and incorrect otherwise.

### v2

@@ 1 Introduction
  • We show that retrieval augmented generation (RAG) significantly improves performance compared with static instruction prompting, as confirmed by extensive benchmarking with top-tier large language models (LLMs).
  • We share our dataset, code, and prompts [^1] https://github.com/Bayer-Group/text-to-sql-epi-ehr-naacl2024 to foster reproducibility and catalyse a community-driven effort towards advancing this research area.
@@ 2 The Dataset
# of question/SQL pairs (all) 306
# of different tables used (all) 13
# of different columns used (all) 44
# logical conditions/query 6.4 (6.7)
# nesting levels/query 1.5 (1.1)
# tables/query 2.7 (0.9)
# columns/query 6.3 (4.7)
# medical entities/query 2.0 (4.1)
Question length [char]/query 91.7 (81.2)
SQL query length [char]/query 796.4 (448.5)
Table 1: Summary statistics of the dataset. For sample statistics, average and standard deviation (in brackets) are reported.
@@ 4.1 Experimental setup
Database and evaluation. The evaluation data reported are obtained querying the DE-SynPUF dataset SynPUF (2010), which is a synthetic dataset that emulates the structure of actual claims data. It includes 6.8 million beneficiary records, 112 million claims records, and 111 million prescription drug events records Gonzales et al. (2023). The same analysis could be applied to any database conforming to the OMOP-CDM, thus potentially allowing access to 2.1 billion patient records Reich et al. (2024). For evaluation, we manually developed a dataset of question-SQL pairs, as detailed in Sec. 2. These are then executed against the DE-SynPUF dataset, and the retrieved data from both reference and generated queries are compared to assess performance. This process reflects the practical use of SQL queries on healthcare databases. To ensure a realistic evaluation setup, the actual question being evaluated is removed from the RAG procedure. A generated query is marked as correct if it retrieves data enabling an answer that aligns with the reference query’s answer (within a 10% tolerance), and incorrect otherwise. The tolerance compensates for variations from GPT-4 Turbo-based medical coding, maintaining the focus on text-to-SQL evaluation accuracy.
@@ 4.2 Experimental results
5 Related Work
Text-to-SQL datasets for EHRs. The MIMIC-SQL dataset  Wang et al. (2020) comprises 10 000 template-generated questions for the MIMIC-III Johnson et al. (2016) database. It contains both question designed to retrieve patient-specific information, and questions on patients counts with logical and basic mathematical operations. Tarbell et al. (2023) noted limited diversity in MIMIC-SQL’s queries, possibly affecting its utility for testing text-to-SQL model generalizability. emrKBQA  Raghavan et al. (2021) contains 1 million patient-specific questions, also based on MIMIC-III. EHRSQLLee et al. (2022) is a dataset created by extracting templates from clinical questions posed by hospital staff, which are then used to generate a comprehensive set of queries for MIMIC-III and eICU Pollard et al. (2018). It relies on an earlier, less performing text-to-text model for query generation  Raffel et al. (2020). All these datasets do not adhere with OMOP-CDM, and they opt for direct string matching for concept retrieval. The closest dataset to ours is the OMOP query library OHDSI (2019); OMOP-CDM-Query-Library (2019), which is a collection of queries in OMOP-CDM. We adapted and included fifteen SQL queries from this library pertinent to epidemiological research into our dataset. Park et al. (2023) use rule-based methods and GPT-4 to translate clinical trial eligibility criteria into SQL queries for OMOP-CDM.
Text-to-SQL with LLMs and in-domain demonstrations. Prompting LLMs has proven effective, often outperforming specialized fine-tuned models in text-to-SQL task Pourreza and Rafiei (2023). Both in-domain Chang and Fosler-Lussier (2023a) and out-of-domain Chang and Fosler-Lussier (2023b) demonstrations improve LLMs’ performance. Gao et al. (2023) explores retrieval scenarios for in-domain demonstration selection. To the best of our knowledge, the exploration of these text-to-SQL methods within EHR (or biomedical) research has not yet extended to small datasets that are critical for industry applications.
6 Conclusion
In this work, we presented the task of answering epidemiological questions using RWD. We demonstrated that RAG is effective in improving performance on all tested scenarios. Our study extends the demonstrated efficacy of RAG from general text-to-SQL benchmarks Gao et al. (2023); Chang and Fosler-Lussier (2023b) to include to small, domain-specific biomedical datasets, underlining its utility in data-scarce industry settings. The primary limitation is the dataset’s limited size and specialized focus on epidemiological questions, suggesting further research should broaden its scope and scale.
7 Acknowledgments
We would like to acknowledge Zuzanna Krawczyk-Borysiak for creating 27 text-SQL pairs; the remaining text-SQL pairs within the dataset were curated by A.Z. We also thank Jared Worful, Melanie Hackl, and Zuzanna Krawczyk-Borysiak for helpful discussions.